OpenStudy (anonymous):

Solve the equation or Inequality. (3x+2)/4 = 9/4 - (3-2x)/6

OpenStudy (anonymous):

\[\frac{3x+2}{4}=\frac{9}{4}-\frac{3-2x}{6}\]?

OpenStudy (anonymous):

yes ive been trying to figure out how to do this

OpenStudy (anonymous):

stupid fractions are getting in the way multiply both sides by \(12\) and start with \[3(3x+2)=27-2(3-2x)\]

OpenStudy (anonymous):

why 12? is that because its a common denominator?

OpenStudy (anonymous):

because it works it works because 12 is the least common multiply of 4 and 6, so if you multiply by 12 you get rid of the denominators

OpenStudy (anonymous):

"least common multiple" i mean

OpenStudy (anonymous):

ah okay thanks

OpenStudy (anonymous):

yw